OpenStudy (anonymous):

A school supply closet contains a box of 21 AA batteries. After testing the batteries, only 10 of them worked. What is the probability of selecting a battery that works followed by a battery that doesn't work without replacing the first? A: 110/441 B: 11/42 C: 3/14 D: 10/49

OpenStudy (anonymous):

@Xmoses1

OpenStudy (anonymous):

@Megan_xoxo1806

OpenStudy (anonymous):

medals for best answer

OpenStudy (anonymous):

Part 1:\[\frac{ 10 }{ 21 }\]Part 2:\[\frac{ 11 }{ 20 }\]---------------\[\frac{ 10 }{ 21 }\times \frac{ 11 }{ 20 }=\frac{ 11 }{ 42 }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome.